OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4). y = four thirdsx + 8 y = four thirdsx + 3 y = -three fourthsx + 8 y = -three fourthsx + 3

OpenStudy (ashes.boom):

if two lines are parallel it means they have the same slope :)

OpenStudy (anonymous):

Can I get help though

OpenStudy (ashes.boom):

Write the equation of the line parallel to 4x - 3y = 12 through the point (-3, 4). 4x - 3y = 12 -4x -4x -3y = -4x + 12 /-3 /-3 y = 4/3x - 4 m = 4/3 (-3, 4) point slope form can help you find the y intercept y - y1 = m(x - x1) y - 4 = 4/3(x - (-3)) y - 4 = 4/3 (x + 3) to find the y intercept just replace x with a zero y - 4 = 4/3 (0 + 3) y - 4 = 4/3 * 3/1 y - 4 = 12/3 or 4 y - 4 = 4 + 4 + 4 y = 8 so thats your y intercept

OpenStudy (ashes.boom):

Does that make sense?

OpenStudy (anonymous):

thank you.

OpenStudy (ashes.boom):

Of course! Let me know if you need help on anything else :)

OpenStudy (anonymous):

\[\huge \color {Green} {Welcome~to}\]\[\huge \color {lightblue} {Open} \color {lightgreen} {Study}\]